Periodic Inventory Using FIFO, LIFO, and Weighted Average Cost Methods The units of an item available for sale during the year were as follows: Jan. 1 Inventory 4 units at $43 Aug. 7 Purchase 18 units at $44 Dec. 11 Purchase 15 units at $45 $172 792 675 $1,639 37 units ermine the inventory cost There are 18 units of the item in the physical inventory at December 31. The periodic inventory system is used. Det using (e) the first-in, first-out (FIFO) method; (b) the last-in, first-out (LIFO) method; and (e) the weighted average cost method cost to two decimal places and your final answer to the nearest whole dollar). a. First-in, first-out (FIFO) b. Last-in, first-out (LIFO) c. Weighted average cost PreviousNext Check My Work 2 more Check My Work uses remaining. Email instructor Save and Exit Submit Assignment for Grading All work saved
Answer #1

FIFO is 807$

LIFO is 788$

Weighted average method is 792$

Date

Purchase

Issue

Inventory

Unit

$/unit

$Total

Unit

$/unit

$Total

Unit

$/unit

$total

Jan1

4

43

172

Aug7

18

44

792

4

18

43

44

172

792

964

Dec11

15

45

675

4

18

15

43

44

45

172

792

675

1,639

in this problem given closing inventory that means there was an issue but not given a particular date.

closing inventory is 18 units. so 19 units were issued.

FIFO

in FIFO first in is first out sofa,

opening inventory is the 4 units at 43$ and Aug 7th inventory 15 units at 44$ are issued

4*43+15*44=832

1,639-822=807

807$ is the FIFO

LIFO

Dec 11th inventory15 units at 45$ and Aug 7th inventory 5 units at 44$ are issued

15*45+5*44=851

1639 - 851= 788

Weighted average cast

the total cost of inventory/units

43,44 and 45 are the weights

average weight is 44

total units are 37

37*44= 1628$

now issued 19 units we can calculate the cost of 44.

19*44= 836

1,628-836=792$